OpenStudy (anonymous):

complete the square (4(x-2)/x-3)+3/x=-3/x(x-3)

OpenStudy (anonymous):

the denominator will be x(x-3)

OpenStudy (anonymous):

(4(x-2)/x-3)+3/x=-3/x(x-3 \[\frac{ 4(x-2) }{ x-3 }+\frac{ 3 }{ x }=\frac{ -3 }{ x(x-3) }\]

OpenStudy (anonymous):

\[\frac{ 4x(x-2)+3(x-3) }{ x(x-3)}=\frac{ -3 }{ x(x-3) }\]

OpenStudy (anonymous):

\[\frac{ 4x^{2}-8x+3x-9 }{ x(x-3) }=\frac{ -3 }{ x(x-3) } \]

OpenStudy (anonymous):

now cancelling both the denominator \[4x ^{2}-5x-6=0\]

OpenStudy (anonymous):

(4x+3)(x-2)=0